OpenStudy (anonymous):

Take 4a - 3b + 2c from 2a - 3b + 4c. a)-2a + 2c b)-2a - 2c c)2a + 2c d)2a - 2c

OpenStudy (anonymous):

I don't think any of your answers are correct....

OpenStudy (anonymous):

Just to clarify, do you mean (21-3b+4c) - (4a-3b+2c)?

OpenStudy (anonymous):

* in the first parentheses, I meant to write 2a

OpenStudy (anonymous):

You add/subtract like terms, or the terms that are similar. I think you worded this problem sort of strangely because doing: (2a - 3b + 4c) - (4a - 3b + 2c) = -2a + 6, which isn't a solution given. Doing (4a - eb + 2c) - (2a - 3b + 4c) = 2a-2c. You set the two equations up on top of each other and subtract each term separately. 4a - 3b + 2c Step 1: 4a - 2b = 2a. 2a - 3b + 4c Step 2: -3b - (-3b) = -3b + 3b = 0 ---------------- Step 3: 2c-4c = -2c 2a + 0 - 2c 2a-2c Does that make sense?